# Recursive Deep Inverse Reinforcement Learning

### v4

@@ 1 Introduction
Previous IRL approaches have included maximum-margin approaches (Abbeel and Ng, 2004), and probabilistic approaches such as (Ziebart et al., 2008). In this work, we build on the maximum entropy IRL framework presented previously (Ziebart et al., 2008). In this framework, training consists of two nested loops. The inner loop approximates the optimal control policy for a hypothesized cost function, while the outer loop minimizes a negative log-likelihood cost function (Ziebart et al., 2008), constructed by sampling a full trajectory from the inner loop’s optimal control policy and by using the expert trajectory that is observed from the expert.
@@
Due to this nested structure, training under the maximum entropy deep IRL in an online fashion becomes very challenging since inner and outer loops need long trajectories and large batch sizes to converge. Moreover, recursive optimization strategies such as Extended Kalman Filter (EKF) sequentially minimize a loss function that is a summation of mean square error of observed and estimated states, and mean squared error of the estimated states and their predicted values produced by assumed model dynamics (Humpherys et al., 2012). Hence, EKF could not be naively leveraged to optimize the negative log-likelihood cost function (Ziebart et al., 2008) since the log of summation term could not be optimized sequentially.
@@
To address these challenges, we first derive an upper-bound cost function for the negative log-likelihood function (Finn et al., 2016b; Ziebart et al., 2008). We then propose a recursive optimization algorithm that minimizes this upper bound using expert demonstrations and sampled trajectories from the inner optimal control policy. This approach alleviates the need to optimize the negative log-likelihood cost function only after collecting all trajectories from the inner-loop policy and the expert. Instead, it enables incremental optimization, processing each expert observation as it arrives. One of the benefits of this approach is that the inner policies generating samples from the partition function (Finn et al., 2016b; Ziebart et al., 2008) are updated on the fly, making the optimization procedure online. Moreover, our proposed approach benefits from a faster convergence speed, which is suitable for online learning scenarios.
@@
Our contributions are twofold: (1) an upper-bound optimization function enabling recursive optimization of maximum entropy cost functions (Ziebart et al., 2008; Finn et al., 2016b), and (2) a deep max-entropy online IRL algorithm, Recursive Deep Inverse Reinforcement Learning (RDIRL), that learns nonlinear cost and reward functions parameterized by neural networks from expert demonstrations as they arrive. Since the inner control policy updates upon each new expert demonstration, our method enables online learning of adversarial agent policies—unlike previous deep learning approaches. We validate our approach on simulated benchmark tasks, where it outperforms leading IRL methods.
@@
IRL, also known as IOC (Finn et al., 2016b), aims to learn reward or cost functions from expert agents operating under optimal control or reinforcement learning policies. Several IOC methods have been developed to recover finite-horizon optimal control cost functions, including approaches based on Karush-Kuhn-Tucker (KKT) conditions(Zhang et al., 2019b, a; Puydupin-Jamin et al., 2012), Pontryagin’s minimum principle (Molloy et al., 2022, 2020; Jin et al., 2020), and the Hamilton-Jacobi-Bellman equation (Pauwels et al., 2014; Hatz et al., 2012).
@@
These methods typically follow a two-stage process: first, a feedback gain matrix is computed from state and control sequences using system identification techniques, and second, linear matrix inequalities are solved to recover the objective-function parameters from the feedback gain matrix. Online variations of IOC methods based on the Hamilton-Jacobi-Bellman equation (Zhao and Molloy, 2024; Molloy et al., 2018, 2020; Self et al., 2020c, a, b) have also been developed. However, both offline and online versions of these methods are generally limited to simple parameter estimation, assume partial knowledge of the expert’s cost function, and do not incorporate deep neural network (Deep Neural Network (DNN)) representations of cost functions.
@@
IRL approaches have also been proposed based on maximum margin (Abbeel and Ng, 2004; Ratliff et al., 2006) and maximum entropy (Ziebart et al., 2008; Boularias et al., 2011). Among these, maximum entropy IRL, as introduced by (Ziebart et al., 2008), has become one of the leading approaches. In this framework, optimization seeks to find reward or cost function parameters that maximize the likelihood of the observed expert trajectory under a maximum entropy distribution. This involves estimating a partition function from samples drawn from a background distribution that represents a control policy (Finn et al., 2016a; Fu et al., 2017), which is dependent on a parameterized cost function. The control policy may range from reinforcement learning (Ho and Ermon, 2016; Fu et al., 2017) to receding horizon optimal control (Xu et al., 2022).
@@
Building on maximum entropy IRL, feature-based methods (Hadfield-Menell et al., 2016; Wu et al., 2020) model the reward function as an inner product between a feature vector $f$ and a parameter vector $\theta$. These methods have been successfully implemented, with the feature characteristics and parameter vector size typically chosen to match the true cost function structure. However, they assume some structural knowledge of the expert’s cost function or domain knowledge (Finn et al., 2016b). Online versions of feature-based maximum entropy IRL have also been developed (Rhinehart and Kitani, 2018; Arora et al., 2021), but they have not yet been extended to include a DNN parameterization of the reward and cost functions.
@@
Similarly, maximum entropy IRL with deep learning representations of the reward function has been successfully implemented (Wulfmeier et al., 2015). These methods, which leverage DNNs for complex reward functions, have gained popularity and become widely used (Finn et al., 2016b; Wulfmeier et al., 2015; Ho and Ermon, 2016; Xu et al., 2019, 2022; Fu et al., 2017, 2019; Yu et al., 2019). As a result, they have emerged as leading IRL approaches, outperforming feature-based methods (Finn et al., 2016b; Xu et al., 2022; Ho and Ermon, 2016).
@@
In this work, we propose a new online IRL method based on the maximum entropy framework (Ziebart et al., 2008; Ziebart, 2010). Unlike other online approaches (Molloy et al., 2018; Self et al., 2020c, b; Molloy et al., 2020; Rhinehart and Kitani, 2018; Arora et al., 2021), our method allows for cost and reward functions to be parameterized using deep neural networks. Our approach is most similar to the algorithm introduced by (Finn et al., 2016a), which minimizes a negative log likelihood cost function and uses Model Predictive Path Integral Control (MPPI) (Xu et al., 2022) as the inner control policy. However, unlike prior work, we recursively adapt the sampling distribution representing the inner control policy each time an expert demonstration is observed.
@@
To summarize, our proposed method is the first to combine several key features into a single, effective algorithm. It can learn adversarial cost functions online, which is critical for applications such as evasion and pursuit. Additionally, it can learn complex, expressive cost functions, parameterized by deep neural networks, eliminating the need for manual design of cost-functions typically required in recursive methods (Molloy et al., 2018; Zhao and Molloy, 2024; Self et al., 2020c). While some prior methods have demonstrated good performance with online IOC (Zhao and Molloy, 2024; Molloy et al., 2020; Self et al., 2020c) and others have explored deep neural network-based cost functions(Finn et al., 2016a; Fu et al., 2017; Ho and Ermon, 2016), to the best of our knowledge, no previous approach has successfully combined all of these properties.
@@ 3.1 Maximum Entropy Inverse Reinforcement Learning
[…]
This approach is equivalent to minimizing the negative log-likelihood of Equation 1 given below (Finn et al., 2016b):
@@
To represent the cost function $c_{\theta}(\tau)$, IOC or IRL feature-based methods typically use a linear combination of hand-crafted features $f:(u,x)\mapsto f(u,x)$, leading to $c_{\theta}(\tau)=\theta^{T}f(u_{t},x_{t})$ (Abbeel and Ng, 2004). This representation is difficult to apply to more complex domains (Finn et al., 2016b). Recent works have focused on the use of high-dimensional expressive function approximators, representing $c_{\theta}(\tau)$ using neural networks, and outperforming feature-based methods (Finn et al., 2016b; Fu et al., 2017; Ho and Ermon, 2016). In this work, we only leverage neural networks to represent the cost function although, other parameterizations could also be used with our method. In practice, the negative log-likelihood in Equation (2) is minimized using gradient descent and batch training. Previous algorithms using deep networks as the cost function parameterization required long and multiple expert demonstrations and sampled trajectories from background policies in order to converge through multiple training iterations. Moreover, training could not proceed before generating all expert and sampled trajectories which restricted it to offline training paradigms. In this work, we introduce a recursive optimization algorithm that adapts network parameters $\theta$ on the fly whenever an expert demonstration is observed.
@@ 3.2 Kalman Filtering
The Kalman Filter (KF) is among the most widely used state estimators in engineering applications. This algorithm recursively estimates the state variables, for example, the position and velocity of a projectile in a noisy linear dynamical system (Lipton et al., 1998), by minimizing the mean-squared estimation error of the current state, as noisy measurements are received and as the system evolves in time (Humpherys et al., 2012). Each update provides the latest unbiased estimate of the system variables. Since the updating process is fairly general and relatively easy to compute, the KF can often be implemented in real-time. When dealing with nonlinear systems extensions of the KF exist such as the EKF which resorts to linearizations using first-order Taylor’s expansions Särkkä and Svensson (2023).
@@
where $X_{n}=\{x_{1},\dots,x_{n}\}$ and $x_{n}$ represents the state of interest at time $n$. Moreover, $Y_{n}=\{y_{1},\dots,y_{n}\}$ where $y_{n}$ represents the measurement data at time $n$. $j_{k}$ represents the cost at time $k$ associated with $x_{k}$ and $y_{k}$, while $J_{n}$ is the cumulative value of $j_{k}$ and represents the cumulative cost associated with trajectories $X_{n}$ and $Y_{n}$. The EKF estimates the state $x_{n}$ that minimizes (3) at time $n$ using second-order Newton method as new measurement $y_{n}$ arrives. Thus, Equation (3) can be re-written as:
@@
The EKF finds $x_{n}$ that minimizes (4) given previous loss function $J_{n-1}$, previous state estimates of $X_{n-1}$, previous measurements $Y_{n-1}$ and current measurement $y_{n}$. In classical Kalman filtering applications such as navigation and target tracking (Ward et al., 2006; Roumeliotis and Bekey, 2000), the goal is to estimate states $x_{n}$ given sequences of noisy (often Gaussian) data $y_{n}$. In this work, however, we aim at estimating the parameters $\theta$ of the cost function $c_{\theta}(\tau)$ from expert demonstration $\tau\in D_{\mathrm{demo}}$ recursively. Inspired by the Kalman filter’s sequential optimization approach described in (Humpherys et al., 2012), we develop a sequential optimization approach to find $\theta$ that maximizes the entropy $p(\tau)$.
4 Upper Bound of the Negative Log-likelihood
To find $\theta$ that maximizes Equation (1), previous approaches (Finn et al., 2016b; Fu et al., 2017; Ho and Ermon, 2016) minimize the negative log-likelihood loss function described in Equation (2). To obtain an online optimizer of (2), we seek to minimize it recursively in a similar fashion to the Kalman filtering optimization described in the previous section. Nevertheless, since recursive minimization requires the loss function to be a summation in a similar fashion to Equations (3) and (4), the second term of Equation (2) which is a log of summation, prohibits recursive optimization. For this reason, we derive an upper bound of Equation (2) which transforms the latter into a summation and then proceed to minimize the upper bound recursively.
In (Matkovic and Pecaric, 2007), authors present a general variant of Jensen’s inequality for convex functions as follows. Let $[a,b]$ be an interval in $\mathbb{R}$, $y_{1},\ldots,y_{N}\in[a,b]$, and $p_{1},\ldots,p_{N}$ be positive real numbers such that $\sum_{n=1}^{N}p_{n}=1$. If $f:[a,b]\rightarrow\mathbb{R}$ is convex on $[a,b]$, then:
@@ 4 Upper Bound of the Negative Log-likelihood
In Matkovic and Pecaric (2007) the authors replaced the function $f$ by the negative $\log$ function, $f=-\log$ which is a convex function. Then Equation 5 can be re-written as:
@@
In what follows, we will consider $N=M$ in (2) for the sake of compactness. Let’s define $p_{n}$ and $y_{n}$ as follows :
@@
where $\tau_{\mathrm{samp}}$ is a trajectory sampled from $D_{\mathrm{samp}}$ and let $y_{n}$ be defined over an interval $[a,b]\in\mathbb{R}$. By replacing (7) in (6) we get the following inequality:
@@
where $K=log(a)+log(b)-2log\left(\frac{a+b}{2}\right)$. Replacing (8) in (2) we can derive an upper bound of (2) as follows:
@@
where $C=\log q(\tau_{i}^{\mathrm{samp}})+K$ and $\tau_{\mathrm{demo}}$ is a trajectory sampled from $D_{\mathrm{demo}}$ representing expert’s trajectory. Since $C$ and $N$ are independent from model parameters $\theta$, minimizing the upper bound of (2) is now equivalent to minimizing the following loss:
$\mathcal{L}_{\mathrm{UB}}=\sum_{i=1}^{N}\left[c_{\theta}(\tau_{i}^{\mathrm{demo}})-c_{\theta}(\tau_{i}^{\mathrm{samp}})\right].$ (10)

### v5

@@ 1 Introduction
Previous IRL approaches have included maximum-margin approaches (Abbeel & Ng, 2004), and probabilistic approaches such as (Ziebart et al., 2008). In this work, we build on the maximum entropy IRL framework presented previously (Ziebart et al., 2008). In this framework, training consists of two nested loops. The inner loop approximates the optimal control policy for a hypothesized cost function, while the outer loop minimizes a negative log-likelihood cost function (Ziebart et al., 2008), constructed by sampling a full trajectory from the inner loop’s optimal control policy and by using the expert trajectory that is observed from the expert.
Due to this nested structure, training under the maximum entropy deep IRL in an online fashion becomes very challenging since inner and outer loops need long trajectories and large batch sizes to converge. Available IRL approaches exploit the fact that it is often feasible to store and process entire state and control sequences in batches(Molloy et al., 2018). In real-time settings with memory, latency and compute constraints, this is generally not feasible.
@@
Recursive optimization strategies such as Extended Kalman Filter (EKF) sequentially minimize a loss function that is a summation of mean square error of observed and estimated states, and mean squared error of the estimated states and their predicted values produced by assumed model dynamics (Humpherys et al., 2012). Hence, EKF cannot be naively leveraged to optimize the negative log-likelihood function (Ziebart et al., 2008) since the log of summation term could not be optimized sequentially. Recent works have proposed moment-matching approaches (Swamy et al., 2021; Zeng et al., 2022; 2025), leading to objective functions that have a simple summation form, making them more suitable for online adaptive learning. However, they are not explicitly derived from maximum entropy IRL, and prior formulations have not been optimized in an online setting.
@@
To overcome this limitation, we require a reformulation of the maximum entropy objective into a structure amenable to recursive optimization. To address this gap, we show that the moment matching loss function introduced in (Swamy et al., 2021) provides an upper bound for the negative log-likelihood objective of maximum entropy IRL (Finn et al., 2016b; Ziebart et al., 2008). We then propose a recursive optimization algorithm that minimizes the moment matching loss using expert demonstrations and sampled trajectories from the inner optimal control policy. This approach alleviates the need to optimize the negative log-likelihood cost function only after collecting all trajectories from the inner-loop policy and the expert. Instead, it enables incremental optimization, processing each expert observation as it arrives.
@@
The main contribution of this work is a deep maximum entropy online IRL algorithm, Recursive Deep Inverse Reinforcement Learning (RDIRL), that learns nonlinear cost and reward functions parameterized by neural networks directly from expert demonstrations as they arrive. Unlike previous deep learning approaches, our method updates the inner control policy after each new expert sample, enabling online adaptation of policies. By processing state–action pairs sequentially, without storing or batching entire trajectories, RDIRL is well-suited for real-time applications with memory and latency constraints. Moreover, because the policy and cost updates occur incrementally, our approach converges significantly faster than competing IRL methods. We validate our approach in simulated benchmark tasks, demonstrating that it outperforms leading IRL methods.
@@
IRL, also known as IOC (Finn et al., 2016b), aims to learn reward or cost functions from expert agents operating under optimal control or reinforcement learning policies. Several IOC methods have been developed to recover finite-horizon optimal control cost functions, including approaches based on Karush-Kuhn-Tucker (KKT) conditions(Zhang et al., 2019b; a; Puydupin-Jamin et al., 2012), Pontryagin’s minimum principle (Molloy et al., 2022; 2020; Jin et al., 2020), and the Hamilton-Jacobi-Bellman equation (Pauwels et al., 2014; Hatz et al., 2012).
@@
These methods typically follow a two-stage process: first, a feedback gain matrix is computed from state and control sequences using system identification techniques, and second, linear matrix inequalities are solved to recover the objective-function parameters from the feedback gain matrix. Online variations of IOC methods based on the Hamilton-Jacobi-Bellman equation (Zhao & Molloy, 2024; Molloy et al., 2018; 2020; Self et al., 2020c; a; b) have also been developed. However, both offline and online versions of these methods are generally limited to simple parameter estimation, assume partial knowledge of the expert’s cost function, and do not incorporate deep neural network (Deep Neural Network (DNN)) representations of cost functions.
@@
IRL approaches have also been proposed based on maximum margin (Abbeel & Ng, 2004; Ratliff et al., 2006) and maximum entropy (Ziebart et al., 2008; Boularias et al., 2011). Among these, maximum entropy IRL, as introduced by (Ziebart et al., 2008), has become one of the leading approaches. In this framework, optimization seeks to find reward or cost function parameters that maximize the likelihood of the observed expert trajectory under a maximum entropy distribution. This involves estimating a partition function from samples drawn from a background distribution that represents a control policy (Finn et al., 2016a; Fu et al., 2017), which is dependent on a parameterized cost function. The control policy may range from reinforcement learning (Ho & Ermon, 2016; Fu et al., 2017) to receding horizon optimal control (Xu et al., 2022).
@@
Building on maximum entropy IRL, feature-based methods (Hadfield-Menell et al., 2016; Wu et al., 2020) model the reward function as an inner product between a feature vector $f$ and a parameter vector $\theta$. These methods have been successfully implemented, with the feature characteristics and parameter vector size typically chosen to match the true cost function structure. However, they assume some structural knowledge of the expert’s cost function or domain knowledge (Finn et al., 2016b). Online versions of feature-based maximum entropy IRL have also been developed (Rhinehart & Kitani, 2018; Arora et al., 2021), but they have not yet been extended to include a DNN parameterization of the reward and cost functions.
@@
Similarly, maximum entropy IRL with deep learning representations of the reward function has been successfully implemented (Wulfmeier et al., 2015). These methods, which leverage DNNs for complex reward functions, have gained popularity and become widely used (Finn et al., 2016b; Wulfmeier et al., 2015; Ho & Ermon, 2016; Xu et al., 2019; 2022; Fu et al., 2017; 2019; Yu et al., 2019). As a result, they have emerged as leading IRL approaches, outperforming feature-based methods (Finn et al., 2016b; Xu et al., 2022; Ho & Ermon, 2016).
@@
In this work, we propose a new online IRL method based on the maximum entropy framework (Ziebart et al., 2008; Ziebart, 2010). Unlike other online approaches (Molloy et al., 2018; Self et al., 2020c; b; Molloy et al., 2020; Rhinehart & Kitani, 2018; Arora et al., 2021), the proposed methodology allows the cost and reward functions to be parameterized using deep neural networks. Our approach is mostly related to the algorithm introduced by (Finn et al., 2016a), which minimizes a negative log-likelihood function and uses Model Predictive Path Integral Control (MPPI) (Xu et al., 2022) as the inner control policy. However, unlike prior work, we recursively adapt the sampling distribution representing the inner control policy each time an expert demonstration is observed.
@@
To summarize, our proposed method is the first to combine several key features into a single effective algorithm. It can learn adversarial cost functions online, which is critical for applications such as evasion and pursuit. Additionally, it can learn complex, expressive cost functions, parameterized by deep neural networks, eliminating the need for manual design of cost-functions typically required in recursive methods (Molloy et al., 2018; Zhao & Molloy, 2024; Self et al., 2020c). While some prior methods have demonstrated good performance with online IOC (Zhao & Molloy, 2024; Molloy et al., 2020; Self et al., 2020c) and deep neural network-based cost functions(Finn et al., 2016a; Fu et al., 2017; Ho & Ermon, 2016; Zeng et al., 2022; Swamy et al., 2021), to the best of our knowledge, no previous approach has successfully combined these two properties.
@@ 3.1 Maximum Entropy Inverse Reinforcement Learning
The partition function $\mathcal{Z}$ is difficult to compute for large or continuous domains, and presents the main computational challenge in maximum entropy IRL. In the sample-based approach to maximum entropy IRL (Finn et al., 2016b; Fu et al., 2017; Ho & Ermon, 2016; Finn et al., 2016a) the partition function $\mathcal{Z}=\int\exp(-c_{\theta}(\tau))d\tau$ is estimated from a background distribution $q(\tau)$ representing the inner control policy, where $\tau$ are sampled from the policy $q(\tau)$. The central idea behind the maximum entropy approach is to estimate $\theta$ that maximizes the likelihood of the entropy cost distribution $p(\tau)$:
$\hat{\theta}=\underset{\theta}{\arg\max}\quad p(\tau).$
This approach is equivalent to minimizing the negative log-likelihood of Equation˜1 given below (Finn et al., 2016b):
@@
To represent the cost function $c_{\theta}(\tau)$, IOC or IRL feature-based methods typically use a linear combination of hand-crafted features $f:(u,x)\mapsto f(u,x)$, leading to $c_{\theta}(\tau)=\theta^{T}f(u_{t},x_{t})$ (Abbeel & Ng, 2004). This representation is difficult to apply to more complex domains (Finn et al., 2016b). Recent works have focused on the use of high-dimensional expressive function approximators, representing $c_{\theta}(\tau)$ using neural networks, and outperforming feature-based methods (Finn et al., 2016b; Fu et al., 2017; Ho & Ermon, 2016). In this work, we only leverage neural networks to represent the cost function although, other parameterizations could also be used with our method. In practice, the negative log-likelihood in equation 2 is minimized using gradient descent and batch training. Previous algorithms using deep networks as the cost function parameterization required long and multiple expert demonstrations and sampled trajectories from background policies in order to converge through multiple training iterations. Moreover, training could not proceed before generating all expert and sampled trajectories which restricted it to offline training paradigms. In this work, we introduce a recursive optimization algorithm that adapts network parameters $\theta$ on the fly whenever an expert demonstration is observed.
@@ 3.2 Kalman Filtering
The Kalman Filter (KF) is among the most widely used state estimators in engineering applications. This algorithm recursively estimates the state variables, for example, the position and velocity of a projectile in a noisy linear dynamical system (Lipton et al., 1998), by minimizing the mean-squared estimation error of the current state, as noisy measurements are received and as the system evolves in time (Humpherys et al., 2012). Each update provides the latest unbiased estimate of the system variables. Since the updating process is fairly general and relatively easy to compute, the KF can often be implemented in real-time. When dealing with nonlinear systems extensions of the KF exist such as the EKF which resorts to linearizations using first-order Taylor’s expansions Särkkä & Svensson (2023).
@@
where $X_{n}=\{x_{1},\dots,x_{n}\}$ and $x_{n}$ represents the state of interest at time $n$. Moreover, $Y_{n}=\{y_{1},\dots,y_{n}\}$ where $y_{n}$ represents the measurement data at time $n$. $j_{k}$ represents the cost at time $k$ associated with $x_{k}$ and $y_{k}$, while $J_{n}$ is the cumulative value of $j_{k}$ and represents the cumulative cost associated with trajectories $X_{n}$ and $Y_{n}$. The EKF estimates the state $x_{n}$ that minimizes equation 3 at time $n$ using second-order Newton method as new measurement $y_{n}$ arrives. Thus, equation 3 can be re-written as:
@@
The EKF finds $x_{n}$ that minimizes equation 4 given previous loss function $J_{n-1}$, previous state estimates of $X_{n-1}$, previous measurements $Y_{n-1}$ and current measurement $y_{n}$. In classical Kalman filtering applications such as navigation and target tracking (Ward et al., 2006; Roumeliotis & Bekey, 2000), the goal is to estimate states $x_{n}$ given sequences of noisy (often Gaussian) data $y_{n}$. In this work, however, we aim at estimating the parameters $\theta$ of the cost function $c_{\theta}(\tau)$ from expert demonstration $\tau\in D_{\mathrm{demo}}$ recursively. Inspired by the Kalman filter’s sequential optimization approach described in (Humpherys et al., 2012), we develop a sequential optimization approach to find $\theta$ that maximizes the entropy $p(\tau)$.
4 Moment Matching as Upper Bound of the Negative Log-likelihood
In this section, we derive an upper-bound of the negative log-likelihood, leading to an optimization problem that is suitable for KF-like online estimation of the parameter vector $\theta$. That is, the resulting upper bound can be written following the same summation structure of equations 3 and 4. The log-sum term in equation 2 prevents direct recursive minimization, but the derived upper bound resolves this issue and enables sequential optimization.
In (Matkovic & Pecaric, 2007), the authors present a general variant of Jensen’s inequality for convex functions as follows. Let $[a,b]$ be an interval in $\mathbb{R}$, $y_{1},\ldots,y_{N}\in[a,b]$, and $p_{1},\ldots,p_{N}$ be positive real numbers such that $\sum_{n=1}^{N}p_{n}=1$. If $f:[a,b]\rightarrow\mathbb{R}$ is convex on $[a,b]$, then:
@@ 4 Moment Matching as Upper Bound of the Negative Log-likelihood
Replacing the function $f$ by the negative $\log$ function, $f=-\log$ which is a convex function, equation 5 can be re-written as Matkovic & Pecaric (2007) :
@@
In what follows, we will consider $N=M$ in equation 2 for the sake of compactness. Let’s define $p_{n}$ and $y_{n}$ as follows :
@@
where $\tau_{\mathrm{samp}}$ is a trajectory sampled from $D_{\mathrm{samp}}$ and let $y_{n}$ be defined over an interval $a,b]\in\mathbb{R}$. By replacing equation 7 in equation 6 we get the following inequality:
@@
where $K=log(a)+log(b)-2log\left(\frac{a+b}{2}\right)$. Replacing equation 8 in equation 2 we can derive an upper bound of equation 2 as follows:
@@
where $C=\log q(\tau_{i}^{\mathrm{samp}})+K$ and $\tau_{\mathrm{demo}}$ is a trajectory sampled from $D_{\mathrm{demo}}$ representing expert’s trajectory. Since $C$ and $N$ are independent from model parameters $\theta$, minimizing the upper bound of equation 2 is now equivalent to minimizing the following loss:
$\mathcal{L}_{\mathrm{UB-MM}}=\sum_{i=1}^{N}\left[c_{\theta}(\tau_{i}^{\mathrm{demo}})-c_{\theta}(\tau_{i}^{\mathrm{samp}})\right].$ (10)
This upper bound has a particularly important consequence: it transforms the maximum entropy IRL objective into a moment-matching loss. This structure is equivalent to recent moment matching formulations in IRL (Swamy et al., 2021; Zeng et al., 2022; 2025), which replace the log-partition function of MaxEnt IRL with expectation-matching objectives between expert and policy distributions. Our derivation shows that moment matching losses, particularly the formulation in (Swamy et al., 2021), can be interpreted as an upper bound of the maximum entropy negative log-likelihood.